OpenStudy (anonymous):

cuberoot(x-2) +7 =28

OpenStudy (anonymous):

\[\sqrt[3]{x-2} +7=28\]

OpenStudy (anonymous):

First you must get the radical by itself

OpenStudy (anonymous):

soo \[\sqrt[3]{x-2}\] = 21

OpenStudy (anonymous):

precisely, then the oppoite of cbrt-ing is ....?

OpenStudy (anonymous):

would you cube it or square it?

OpenStudy (anonymous):

cube it, since we're dealing with cbrt, rememebr like goes with like

OpenStudy (anonymous):

21 cubed is 9,261

OpenStudy (anonymous):

so x-2 = 9,216

OpenStudy (anonymous):

then get x by itself

OpenStudy (anonymous):

add 2.. so x=9,218

OpenStudy (anonymous):

Alright now plug it back in to your original equation and see if both sides match up.

OpenStudy (anonymous):

yeah, it does but its not an answer choice

OpenStudy (anonymous):

lol.. is 9263 a choice?

OpenStudy (anonymous):

Sorry, my bad, and thats weird, whata re the choices?